Wednesday, October 13, 2010

Math Studies II 10-13-10

NO BELLRINGER!!
DUE WEDNESDAY 10/20/10
  • INTRODUCTION of the IA

Should include:

  1. Title
  2. Clear statement of the task:
  3. Hypothesis
  4. A plan ( in steps)
  5. If you have a survey or experiment make sure you attach a sample of it to your introduction.

Also DONT USE WORDS SUCH AS

  1. Causation
  2. Random

Linear Correlation

  • Used for quantitative data > get it through R-Value

We worked on pg. 581 of our text book problem number 3

  • Diagnostics on (calculator)
  • L1 = Judge A L2= Judge B
  • Enter Data in calculator
  • Then go to stat, calc, and choose nr. 4=lin.reg
  • Your results should look like this:
  • y=ax+b
  • a=1.297979798
  • b=-2.411616162
  • r2=.8883643358
  • r=.9425308142

Plug everything in the equation of a line formula given = y= 1.30x+-2.41

Then press Stat, plot 1 (on), Fix windows, and Graph it.... Now you have your graph :)

NO HOMEWORK FOR TONIGHT BUT YOU CAN START ON YOUR INTRODUCTION!!!

Question of the day: What is the difference between r2 and r?

Wednesday, October 6, 2010

Math Studies II

Bellringer: Consider the question; Students' math grades and science grades are an indicator of college success, what data could i collect to pursue an evaluation of my hypothesis?

Agenda: I: Bell Ringer
II: Discussion of student topics

By 10.8.10 3:00pm !
Hw- In an e-mail to susysullivan@gmail.com

1. A topic for your project
2. A hypothesis to be explored
3. A minimum of 5 Data points to be collected

BIG PICTURE: Mathematics provides the tools to analyze data relationships

OBJECTIVE: students will find a project of interest to pursue for their internal assessment.


QUESTION OF THE DAY:
If i wanted to find data points on the effect of music on teenagers and their study habbits what could they be?

Monday, October 4, 2010

Math studies II 10-4-10

Agenda
  • Bellringer
  • Math Studies paper

Bellringer: Name 5 different Data points you could collect if your math Studies project was " The Relationship of Gross Domestic Product and Infant Health"

Possible answer

death rate per 1000 live births

Objective:Students will indentify a valid project that can be completed in 7 weeks.

Big picture: Math and statistics help und understanding the world around us.

Notes

  • $4.3 trillion is Japan's Gross Domestic Income
  • Based on population
  • Infant mortality per every 1000 births
  • Deaths per Day
  • By gender/race/ethnicity
  • Family income
  • Goverment support of medicine

Amount of Data

  • About 100 Data Points
  • 4-5 data points

Types of Projects you can do

  • execute an Experiment
  • Execute a Survey
  • Written by yourself or data that has be published and is valid.
  • Liekert Scale
  • from 1-5

Question of the Day: Name 3 data points that can be used to if my project question is "the relationship between obesity and health probelms"??

Friday, October 1, 2010

IB Math 2 10/1/10

Agenda

· Reminders

· Homework Review

· Internal Assessment

!!For All Students that were not at school today (10-1-10) all make up points from the test are due Monday!!

Home Work (We went over EX 18B.2 Questions 1&2)

1. The scatter plot alongside shows the association between the number of cars crashes in which a casually occurred and total number of car crashes in each year from 1972 to 1994. Given that r value is 0.49:

A) Find r^2

R=0.49

So what you do is this .49x.49= r^2=0.240

B) Describe the association between these variables.

Its Weak

2. In an investigation to examine the association between the Tread depth (y mm) and the number of kilometers travelled (x thousand) a sample of 8 tyres of the same brand was taken and the result are given below.

Kilometers (x thousand)

14

17

24

34

35

37

38

39

L1

Tread depth (y mm)

5.7

6.5

4.0

.30

1.9

2.7

1.9

2.3

L2

A) Draw A Scatter plot of the data

To make a scatter plot

Use your calculator and press stat and enter in edit

Place the following numbers in the appropriate Colum as shown by the red marking on top

After you done that go to

2nd Y= and have the first plot on and the rest off

Then go to window and type in

Xmin=10

Xmax=40

Xscl=2

Ymin=0

Ymax=7

Yscl=5

X res remains as is (1)

Then plot

B) R= .951

R^2= .+904

Internal Assessment

!!Due Monday!! 10/04/10

Everyone must bring in a plan for what you will want to do for your math internal assessment. Use the big packet that ms.s gave you! For an idea. You can Google math internal assessment for ideas.

!!Due Thursday!! 10/7/10

Using your plans that you brought on Monday. You must bring in how you are planning to do your internal assessment.

Wednesday, September 29, 2010

Math Studies Dos

Sorry it took me so long to post this. I don't have a problem for guys but if I think of one if i can actually understand what I'm doing I'll post it up. So for now, here are the notes for today.

Agenda:
I. Find Correlation Coefficient
II. Mastering Calculator
III. Your Math Project

Objective: Students will be able to find the correlation coefficient using calculators.

On your calculator go to STAT. Click on No.1: Edit. Type in the numbers below.

L1: 2, 7, 12, 4, 9, 20, 8, 3, 32
L2: 7, 20, 32, 16, 27, 58, 5, 10, 104

Once done exit out pressing 2nd then Mode. Click on STAT then go right for CALC. Click on No. 1: 1-Var Stats, press ENTER again.

X Bar should equal and rounded to 10.78.

Next hit 2nd and ENTER. Now click 2nd and 2 for L2. Press ENTER.

X Bar should equal 31.

Next press STAT then EDIT. Highlight L3. Type in L1(2nd and 1) minus 10.78. L3(1) should now equal -8.78.

Next highlight L4. Type in L2(2nd and 2) minus 31. L4(1) should now equal -24.

Next highlight L5. Type in L3(2nd and 3) times L4(2nd and 4). L5(1) should now equal 210.72.

Clear Screen(2nd and Mode) then click clear.

Now press 2nd then STAT. Go to the right to MATH. Press No. 5 for sum. Sum(L5(2nd and 5)). You should get 2396. Divide that by 9 for the number of data. Press ENTER. You should now get 266.2.

Once again press 2nd then STAT. Go to the right to MATH. Press No. 5 for sum. Sum(L4(2nd and 4) squared). Press ENTER. You should get 8114.

Again press 2nd then STAT. Go to the right to MATH. Press No. 5 for sum. Sum(L3(2nd and 3) squared). Press ENTER. You should get 745.6 round to 746.

Now square root 8114 divided by 9. You should get 30.
Again square root 746 divided by 9. You should get 9.1.

Now multiply 30 and 9.1. The answer should be 273.

R= 266.2 divided by 273. Which equals .975. Square that to get .95 as your answer.

I really don't get this part here so yeah.

Press 2nd and 0. Go down to DiagnosticOn Press ENTER and ENTER again. It should say done.

Go to STAT then go to the right for CALC. Go down to No.4 for LinReg(a+b). Press ENTER and ENTER again.
y=a+b
a=3.213710879
b=-3.636661699
r2=.948983395
r=.9741577875


HOMEWORK IS ON PAGE 582. THE PROBLEMS ARE 1-4.

Monday, September 27, 2010

Math Studies 2

Septmeber 27, 2010

Bell Ringer
John has a 10% chance of solving a math problem.
Kate has a 20% chance of solving a math problem.
Joan has a 30% chance of solving a math problem.

If these chances are independent what is the probability that the problem will be solves if they work together?

Agenda
1. Bell Ringer
2. Correlation of 2 variables.
3. using variables in analysis.

Objective: students will be able to construct and evaluate a 2 variable scatter plot.

Big Picture: Two variable analysis allows a person to see patterns in data that may indicate a relationship.

Question
Shawnelle has a 13% chance of solving an art problem.
Rosanna has a 26% chance of solving an art problem.
Jack has a 90% chance of solving an art problem.
If these chances are independent what is the probability that the problem will be solves if they work together?

Wednesday, September 22, 2010

Review Set 14A

1. Donald keeps record of the number of clients he telephones over a consecutive period of days.
a. for how many days did the survey last?
b.estimate Donald's chances of telephoning:
i) no clients in a day
ii) four or more clients on a day
iii). less than three clients on a day

a. 1+6+12+9+6+3+2=39 total number of days
b. i). 1/39
ii). 6+3+2=11/39
iii). 1+6+12=19/39
2. David conducted a survey to determine the ages of people walking through a shopping mall. the results are shown in your book. find , correct to 3 decimal places the estimated probability that the next person David meets in the shopping mall is:
a. between 20 and 39 years of age
b. less than 40 years of age
c. at least 20 years of age

a. 45/112=0.4017~0.402
b. 45+19=64/112=0.5714~0.571
c. 45+37+11/112 or 112-19/112=93/112=0.8303~0.830
3. A farmer fences his rectangular property into 9 rectangular paddocks:
If a paddock is selected at random,what is the probability that:
a. it has no fences on the boundary of the property
b. it has one fence on the boundary of the property
c. it has two fences on the boundary of the property?

a. 1/9 because all but one is on a boundary
b. 4/9
c. 4/9
4. When a box of drawing pins dropped on to the floor it was observed that 47 landed on their backs and31 landed on their sides.Find correct to 2 decimal places,the estimated probability of a drawing pin landing:
a. on its back
b. on its side


47-Back, 31-Side
78 total number of pins
a. 47/78=0.602~0.60
b. 31/78=0.397~0.40
5. A saw mill receives logs of various lengths from a plantation. the length of a log is important in being able to produce timber of the length required.the following data indicates the lengths of at least100 logs received.
a. what is the probability of a log being less than 11 metres long arriving at the saw mill?
b. What is the probability of a log being longer than 15 metres arriving at the saw mill?
c. in the next batch of 50 logs how many would be expected to be between 11 m and 15 m long?

a. L < 11=3+4+14/100=21/100
b. L > 15=7+8/100=15/100
c. 11< L < 15=12+18+20+14=64/2=32 per 50 logs
6.At peak hour railway crossings are closed 30% of the time. If you have to drive through three railway crossings during peak hour, what are the chances you will have to stop at least once?

o-open=.7, c-closed=.3
Possible combination's
ooo-.7*.7*.7=.343
ooc
occ
coo
coc
cco
ccc
1-P(ooo)
1-.343=0.657
7.In a golf match, Annette has a 70% chance of hitting the green when using a nine iron and Kari has a 90% chance when using the same club. If, at a particular hole,they both elect to use a nine iron to play to the green, determine the probability that:
a. both hit the green
b. neither hits the green
c. at least one hits the green
d. only Anette hits the green

a..7*.9=.63
b. .3*.1=.03
c. 1-P(of no one)
1-.03=.97
d. .7*.1=.07
8. Jar A contains 3 white and 2 red marbles. Jar B contains 6 white and 4 red marbles. A jar is selected at random and then two marbles are selected WITHOUT replacement. Determine the probability that:
a. both marbles are white
b. two red marbles are picked from Jar A

Jar A: 3 whites, 2 red
Jar B: 6 whites, 4 red
a. Jar A- 1/2*3/5*2/4=6/40
Jar B- 1/2*6/10*5/9=30/180
6/40+30/180=19/60 (add because they're independent)
b. 1/2*2/5*1/4=2/40
QUESTION OF THE DAY
If P(C)= .10 P(D)= .3 and P(C U D)= .11 are C and D independent events?